44. One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
A) 5
B) 7
C) 17
D) 34
I really appreciate your help!!!!

Found 2 solutions by checkley71, tutorcecilia:
Answer by checkley71(8403)

THE NUMBERS ARE X & (X+12)
X+2(X+12)=39
X+2X+24=39
3X+24=39
3X=39-24
3X=15
X=15/3
X=5 THUS
X+12=5+12=17 C) 17 IS THE CORRECT ANSWER.

Answer by tutorcecilia(2152)

Translate English into Math:
x= One number
x+12 = the other number is 12 more than another
x+2(x+12) = The sum (addition) of the smaller number and twice the larger
Total= is 39
.
Putting it all together:
x+ 2(x+12)=39
x+2x+24=39
3x+24=39
3x+24-24=39-24
3x=15
x=5
.
The smaller number = 5
The larger number = (x+12)=(5+12)=17
Plugging the values back into the original equation:
x+ 2(x+12)=39
(5)+ 2((5)+12)=39
(5)+ 2((5)+12)=39
5+2(17)=39
5+34=39
39=39 [checks out]